You have 4 gallons of juice that is 20% apple juice. How much apple juice do you have to add to get a mixture that is 30% apple juice?
Answer by ptaylor(2198)

Let x=amount of pure apple juice that needs to be added
Now we know that the amount of pure apple juice that exists before the mixture takes place is equal to the amount of pure apple juice that exists after the mixture takes place.
Pure apple juice before the mixture takes place:
0.20*4+x
Pure apple juice after the mixture takes place:
0.30(4+x)
Soooooo:
0.20*4+x=0.30(4+x) simplify
0.80+x=1.2+0.30x subtract 0.30x and also 0.80 from each side
x-0.30x=1.2-0.80
0.70x=0.40
x=4/7 =0.57 gal--amount of pure apple juice needed
CK
0.20*4+0.57=0.30(4.57)
0.80+0.57=1.371
1.37~~~1.371
Hope this helps---ptaylor